What amount of magnesium nitride is required to react completely with 0.25 mol of water? What amounts of the products are expected?
1
Expert's answer
2020-08-28T06:26:32-0400

Mg3N2 + 6H2O --> 3Mg(OH)2 + 2NH3

According to the equation, the amounts in moles are:

n (Mg3N2) = 0.25 * 1/6 = 0.042 mol

n (Mg(OH)2) = 0.25 * 3/6 = 0.13 mol

n (NH3) = 0.25 * 2/6 = 0.083 mol

The masses are:

m (Mg3N2) = 0.25 * 1/6 * 100.95 = 4.2 g

m (Mg(OH)2) = 0.25 * 3/6 * 58.32 = 7.3 g

m (NH3) = 0.25 * 2/6 * 17.03 = 1.4 g